OpenStudy (anonymous):

if 10^x/10^y=100 then x=

OpenStudy (anonymous):

10 obviously.

OpenStudy (anonymous):

answer choice is y+2,y-2,2-y,2y,y/2

OpenStudy (anonymous):

I think your teacher or whatever is trolling you. But if I had to make it simpler it would be y+2.

OpenStudy (anonymous):

ty

jimthompson5910 (jim_thompson5910):

10^x/10^y=100 10^(x-y)=10^2 x-y = 2 x = y+2

jimthompson5910 (jim_thompson5910):

The idea is to get the bases the same, so you can equate exponents and solve